# DRC: Enhancing Personalized Image Generation via Disentangled Representation Composition

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Personalizationccs: Information systems Multimedia content creation
@@ 1. Introduction
  • LMM-based method, Pigeon (Xu et al., 2024c), leverages LMMs to convert history images and multimodal instructions into visual and textual tokens, and then infer user preferences and requirements for generation, showcasing the superiority over previous methods. Nevertheless, the entangled nature of style and semantic features in images makes it difficult for LMMs to implicitly identify and utilize task-relevant cues to guide generation. As depicted in Figure 1(b), Pigeon fails to accurately capture the desired semantics specified by the reference image (left) or preserve the user-preferred style in history images (right), which we refer to as Guidance Collapse.
@@
We conduct extensive experiments on two scenarios, following (Xu et al., 2024c): personalized sticker and movie poster generation, comparing DRC with various baselines. Quantitative and qualitative evaluations demonstrate that DRC shows competitive performance while achieving effective disentanglement to mitigate the guidance collapse issue. Our code and data are publicly available at https://anonymous.4open.science/r/DRC-0908.
@@
  • We design a two-stage learning strategy that first establishes effective style–semantic disentanglement for LMMs through a reconstruction-driven paradigm with difficulty-aware optimization, and then leverages semantic-preserving augmentations to effectively transfer the disentanglement capability to personalized image generation.
@@ 2. DRC
  • Disentanglement learning, which employs a self-supervised approach to disentangle the semantic content and visual style of each image; and
@@
Figure 3. Illustration of DRC, which consists of two stages: 1) disentanglement learning, aiming to separate the semantic content and visual style of a given image; and 2) personalized modeling, which adapts LMMs to generate personalized images by integrating style patterns from user historically interacted images with the semantic information form the reference image.
@@ 2.1.1. Contrastive Samples
  • Positive-semantic image: We use LLaVA (Liu et al., 2023) to extract textual semantic descriptions of each anchor image, and then employ SDXL (Podell et al., 2024) to synthesize images that retain the same semantics but differ in style.
@@ 2.1.2. Dual-tower Disentangler
We implement $\text{{StyleTower}}(\cdot)$ and $\text{{SemTower}}(\cdot)$ to extract style and semantic representations from input images, respectively. Both towers share the same architecture—comprising a cross-attention layer followed by a feed-forward layer—and operate in a similar manner. We use $\text{{StyleTower}}(\cdot)$ as an example to illustrate their functionality.
For an image triplet $<\bm{x}_{a},\bm{x}_{s},\bm{x}_{m}>$, we focus only on $\bm{x}_{a}$ and $\bm{x}_{s}$ when extracting style features. To extract the style from either image, the StyleTower uses the target image as the query and the other as the reference (serving as the key and value) in the cross-attention layer. Let $\bm{F}_{a}^{\text{sty}}$ and $\bm{F}_{s}^{\text{sty}}$ denote the extracted style representation of $\bm{x}_{a}$ and $\bm{x}_{s}$, respectively. They can be formulated as follows:
@@
where $q$, $k$, and $v$ denote the query, key, and value in the cross-attention. $\bm{E}_{a}$ and $\bm{E}_{s}$ represent the visual token embedding sequence of $\bm{x}_{a}$ and $\bm{x}_{s}$, respectively, obtained via the LMM’s visual tokenizer and embedding lookup. By ensuring that the query and key originate from images with similar style characteristics, the attention mechanism is expected to capture the shared style features. Following a similar process, the SemTower extracts the semantic information from $\bm{x}_{a}$ and $\bm{x}_{m}$, yielding semantic representations $\bm{F}_{a}^{\text{sem}}$ and $\bm{F}_{m}^{\text{sem}}$, respectively.
@@ 2.1.3. Fusion for Reconstruction
After obtaining the disentangled representations, we reconstruct the original images through a recovery process. To align with the LMM generation pipeline, we first employ a fusion network (FusionNetwork) to encode the combination of style and semantic features into compact latent instructions, which are then used to prompt the LMM for image generation. Ideally, combining any style feature from $\{\bm{F}_{a}^{\text{sty}},\bm{F}_{s}^{\text{sty}}\}$ with any semantic feature from $\{\bm{F}_{a}^{\text{sem}},\bm{F}_{m}^{\text{sem}}\}$ could allow for the reconstruction of the original anchor image. Accordingly, all possible combinations are considered for image reconstruction. Using the combination $(\bm{F}_{a}^{\text{sty}},\bm{F}_{m}^{\text{sem}})$ as an example, the fusion process is formulated as follows:
@@ 2.1.4. Training
During training, we enforce the model to reconstruct the original image using any combination from $\mathcal{Z}=\{\bm{z}{(a,a)},\bm{z}{(a,m)},\bm{z}{(s,a)},\bm{z}{(s,m)}\}$. This encourages the StyleTower to extract the shared information—explicitly defined as style during the construction of contrastive samples—between the anchor and positive-style images, and the SemTower to capture the shared information—explicitly defined as semantics—between the anchor and positive-semantic images, thereby promoting meaningful disentanglement.
@@ 2.2. Personalized Modeling
Equipped with the well-trained disentanglement mechanism, DRC facilitates personalized image generation by constructing user-specific latent instructions from the historically interacted images and the given reference image to steer generation using the LMM, as shown in the right part of Figure 3. The historical images provide user-preferred style information, while the reference image supplies the semantic information the user requires. Next, we describe how user-specific latent instructions are constructed for personalized image generation, followed by the design of the training and inference procedures.
@@ 2.2.1. User-specific Latent Instructions
where $\bm{m}_{i}^{\text{sty}}$ and $\bm{m}^{\text{sem}}$ denote the learned masks for $\bm{F}_{i}^{\text{sty}}$ and $\bm{F}^{\text{sem}}$, respectively. Following (Xu et al., 2024c), the masks are generated using a transformer encoder-based mask generator that predicts binary masks. Formally,
@@
$\bullet$ Obtaining Latent Instructions via Style and Semantic Fusion.
Finally, we fuse the obtained user-preferred style representation with the reference semantic representation to form the user-specific latent instruction $\tilde{\bm{z}}$ for the user $u$ using the FusionNetwork:
@@ 3. Experiments
DM-based TI 18.67 40.90 36.58 0.7654 7.63 32.91 53.67 48.50 105.48
LLM-based PMG 19.16 47.34 39.15 0.7383 13.94 18.31 45.45 37.80 84.91
LLaVA* 17.88 47.26 42.59 0.7575 8.11 17.54 42.65 39.25 93.23
LLaVA 18.72 37.44 33.19 0.7552 7.64 27.02 49.15 43.88 95.19
LaVIT* 18.77 53.63 50.96 0.6855 9.94 15.49 40.76 39.09 107.53
LaVIT 16.39 40.56 40.84 0.7377 8.59 25.74 70.80 69.93 83.39
Pigeon 22.03 61.64 57.26 0.6800 6.99 25.74 50.66 48.34 93.60
LMM-based DRC 23.19 65.10 60.00 0.6770 6.52 21.88 48.30 45.79 92.31
Recon 16.30 40.60 40.76 0.7370 8.84 25.84 71.09 70.14 83.57
Reference Grd 16.93 45.00 43.71 0.6443 8.10 28.95 100.00 100.00 -
@@
Methods CS $\uparrow$ CIS $\uparrow$ DIS $\uparrow$ LPIPS $\downarrow$ StyleSim $\downarrow$ CS $\uparrow$ CIS $\uparrow$ DIS $\uparrow$ FID $\downarrow$
DM-based TI 12.41 28.29 19.18 0.7721 10.83 33.84 43.53 39.81 79.77
LLM-based PMG 13.61 25.11 22.73 0.7692 16.69 15.60 27.29 25.15 77.25
LLaVA* 12.24 29.60 19.73 0.7607 10.02 14.55 31.76 21.99 73.77
LLaVA 12.62 30.64 19.33 0.7690 10.67 30.53 48.50 41.45 54.55
LaVIT* 12.64 28.23 17.50 0.7546 8.87 19.39 36.93 37.71 50.08
LaVIT 13.86 30.49 19.95 0.7548 10.97 25.15 46.02 60.07 33.53
Pigeon 15.19 37.42 17.70 0.7496 10.54 27.30 47.79 41.35 41.35
LMM-based DRC 15.24 37.53 19.26 0.7538 9.57 25.71 48.26 43.45 42.20
Recon 13.85 30.33 19.95 0.7548 9.82 25.29 46.08 60.52 33.74
Reference Grd 10.94 51.34 20.75 0.7502 9.89 31.81 100.00 100.00 -
@@ 3.1.2. Baselines
We compare DRC with several DM-based, LLM-based, and LMM-based methods: 1) Textual Inversion (TI) (Gal et al., 2023) models user preferences using a word embedding, which is then used to guide diffusion models (DMs) in generating personalized images. 2) PMG (Shen et al., 2024) models image features in the form of keywords and converts them into embeddings that guide image generation. 3) LLaVA (Liu et al., 2023) is a widely used LMM primarily designed for image understanding. With an external text-to-image generator, it can also be adapted for image generation. 4) LaVIT (Jin et al., 2024) is another LMM that quantizes images into discrete visual tokens using a visual encoder and quantizer. These visual tokens are then processed by a LLM and subsequently used by DMs for image generation. 5) Pigeon (Xu et al., 2024c) builds upon LaVIT by introducing a mask generator to remove noise from images before passing them to the LMM for improved personalized image generation. For fair comparison, both DRC and Pigeon are evaluated under the SFT setting only.
@@ 3.1.3. Evaluation Metrics
  • Style Alignment. To assess style preservation, we measure the visual similarity between generated images and user history images. Specifically, LPIPS (Zhang et al., 2018) and StyleSim (Huang and Belongie, 2017) are used to evaluate perceptual and style-level similarity. We further compute the CLIP Image Score (CIS) and DINO Image Score (DIS) by measuring the cosine similarity of visual features extracted by CLIP (Radford et al., 2021) and DINO (Oquab et al., 2023), respectively. In addition, the CLIP Score (CS) is calculated between the generated images and textual descriptions of the history images.
@@ 3.1.4. Implementation Details
All baselines are trained using their default settings. In DRC, at the disentanglement learning stage, we set the learning rate to $1e^{-4}$. At the personalized modeling stage, following (Xu et al., 2024c), we set the learning rate to $1e^{-5}$ and style mask ratio $\alpha_{s}$ to 0.2. During training, the semantic mask ratio $\alpha_{m}$ is sampled from a uniform distribution, $\alpha_{m}\sim\mathcal{U}(0,1)$. During inference, we select the optimal semantic mask ratio $\alpha_{m}\in\{0.0,0.1,\dots,1.0\}$ for each image based on the F1 score of the style CIS and the semantic CS metrics. We use LoRA (Hu et al., 2022) to finetune our model, with lora r set to 8 and lora $\alpha$ set to 16. All experiments are conducted on a single NVIDIA A100 GPU.
@@ 3.2. Overall Performance (RQ1)
Table 2 presents the performance comparison between DRC and all baseline methods. We conclude the observations as follows:
  • DM-based TI outperforms other baselines in semantic alignment, primarily capturing the semantic content of the reference image while struggling to effectively extract user preferences from the history images. As a result, its outputs often fail to align with user style preferences.
@@
  • For the pre-trained LLaVA and LaVIT, it is challenging to effectively extract the task-relevant information from the mixed inputs of multiple images and textual instructions, resulting in unsatisfactory performance. After fine-tuning, the models tend to focus on reconstructing the reference image while overlooking the style preference embedded in the user-interacted history images, which can be attributed to the lack of high-quality supervised datasets.
  • DRC and Pigeon achieve the best overall performance, excelling in both semantic alignment and personalization. Notably, DRC can achieve comparable, even superior performance to Pigeon in most metrics. The explicit disentanglement strategy allows the model to effectively separate and integrate style and semantic information, resulting in substantial performance gains. More importantly, as illustrated in Figure 4, DRC can accurately extract and fuse user-preferred styles and reference semantics for personalized sticker generation, effectively mitigating the guidance collapse issue observed in Pigeon. For instance, in the first case, DRC successfully captures the owl character style from the user-interacted history images and combines it with the shopping semantics from the reference image to produce a shopping-themed owl sticker. In contrast, Pigeon leans toward reconstructing the reference image, neglecting the personalized style features present in the history. More examples are provided in Appendix A.2.
@@ 3.3.1. Effect of disentanglement learning (RQ2)
Table 3. Effects of the designs in disentanglement learning. “Imp.” denotes importance sampling trick. “Fusion” denotes FusionNetwork. “Attn.” denotes the attention-based disentangler.
@@ 4. Related Work
$\bullet$ Personalized Generation. With the rapid development of large generative models, such as DMs (Podell et al., 2024; Esser et al., 2024), LLMs (Grattafiori et al., 2024; Guo et al., 2025; Yang et al., 2024a), and LMMs (Zhou et al., 2025; Chen et al., 2025; Xie et al., 2025), personalized generation has attracted growing attention across diverse modalities, including text (Salemi et al., 2024; Qiu et al., 2025), image (Xu et al., 2024b, c), 3D (Raj et al., 2023; Liu et al., 2024), video (Wang et al., 2024a; Wei et al., 2024b), audio (Li et al., 2024; Plitsis et al., 2024), and cross-modal generation (Pham et al., 2024; Nguyen et al., 2024). By synthesizing content that aligns with specific user preferences and requirements, personalized generation paves the way for more engaging and user-centric experiences across various application scenarios (Xu et al., 2025).
@@
$\bullet$ Large Multimodal Models. Recently, an increasing number of studies have explored LMMs for visual understanding and generation. Some approaches, such as LLaVA (Liu et al., 2023) and Mini-GPT4 (Zhu et al., 2024), leverage pre-trained vision encoders to extract visual features, which are then projected into LLMs via a dedicated mapping layer for image understanding. To enable both visual understanding and generation, models like LaVIT (Jin et al., 2024) and SEED-X (Ge et al., 2024) introduce dedicated visual tokenizers that convert images into discrete visual tokens through a learnable codebook, supporting autoregressive generation of both textual and visual tokens. In contrast, DreamLLM (Dong et al., 2024) enhances LLMs by introducing a special token to indicate image generation positions, along with a series of queries that encode historical semantics to guide the image generation process within a diffusion model-based decoder. More recently, advanced LMMs with native image generation capabilities, such as Chameleon (Team, 2024), Show-o (Xie et al., 2025), Transfusion (Zhou et al., 2025), and GPT-4o (OpenAI, 2024), have demonstrated impressive performance in multimodal understanding and generation, establishing a solid foundation for future advancements in personalized image generation. Notably, our proposed DRC adopts LaVIT as its backbone, while its design remains compatible with other LMM architectures, which we leave as an exciting direction for future exploration.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755051isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Personalizationccs: Information systems Multimedia content creation
@@ 1. Introduction
  • LMM-based method, Pigeon (Xu et al., 2024c), leverages LMMs to convert history images and multimodal instructions into visual and textual tokens, and then infer user preferences and requirements for generation, showcasing the superiority over previous methods. Nevertheless, the entangled nature of style and semantic features in images makes it difficult to implicitly identify and utilize task-relevant cues to guide generation. As depicted in Figure 1(b), Pigeon fails to accurately capture the desired semantics specified by the reference image (left) or preserve the user-preferred style in history images (right), which we refer to as Guidance Collapse.
@@
We conduct extensive experiments on two scenarios, following (Xu et al., 2024c): personalized sticker and movie poster generation, comparing DRC with various baselines. Quantitative and qualitative evaluations demonstrate that DRC shows competitive performance while achieving effective disentanglement to mitigate the guidance collapse issue. Our code and data are publicly available at https://github.com/ZhengWwwq/DRC.
@@
  • We design a two-stage learning strategy that first establishes effective style–semantic disentanglement for LMMs through a reconstruction-driven paradigm with difficulty-aware optimization, then utilizes semantic-preserving augmentations to transfer the disentanglement capability to personalized image generation.
@@ 2. DRC
  • Disentanglement learning, which employs a self-supervised approach to disentangle the style and semantics of each image.
@@
Figure 3. Illustration of DRC, which consists of two stages: 1) disentanglement learning, aiming to separate the semantic content and visual style of a given image; and 2) personalized modeling, which adapts LMMs to generate personalized images by integrating style patterns from user historically interacted images with the semantic information from the reference image.
@@ 2.1.1. Contrastive Samples
  • Positive-semantic image: We use LLaVA (Liu et al., 2023) to extract textual semantics of each anchor image, then employ SDXL (Podell et al., 2024) to create images that retain the same semantics but differ in style.
@@ 2.1.2. Dual-tower Disentangler
[…]
For an image triplet $<\bm{x}_{a},\bm{x}_{s},\bm{x}_{m}>$, we focus only on $\bm{x}_{a}$ and $\bm{x}_{s}$ when extracting style features. To extract the style from either image, the StyleTower uses the target image as the query and the other as the key and value in the cross-attention layer. Let $\bm{F}_{a}^{\text{sty}}$ and $\bm{F}_{s}^{\text{sty}}$ denote the extracted style representation of $\bm{x}_{a}$ and $\bm{x}_{s}$, respectively. They can be formulated as follows:
@@
where $q$, $k$, and $v$ denote the query, key, and value in the cross-attention. $\bm{E}_{a}$ and $\bm{E}_{s}$ represent the visual token embedding sequence of $\bm{x}_{a}$ and $\bm{x}_{s}$, obtained via the LMM’s visual tokenizer and embedding lookup. Since the query and key originate from images with similar style characteristics, the attention mechanism is expected to capture the shared style features. Following a similar process, the SemTower extracts the semantic information from $\bm{x}_{a}$ and $\bm{x}_{m}$, yielding semantic representations $\bm{F}_{a}^{\text{sem}}$ and $\bm{F}_{m}^{\text{sem}}$, respectively.
@@ 2.1.3. Fusion for Reconstruction
With the disentangled representations, we reconstruct the original images through a recovery process. To align with the LMM generation pipeline, we first employ a fusion network to encode the combination of style and semantic features into compact latent instructions, which are then used to prompt the LMM for image generation. Ideally, combining any style feature from $\{\bm{F}_{a}^{\text{sty}},\bm{F}_{s}^{\text{sty}}\}$ with any semantic feature from $\{\bm{F}_{a}^{\text{sem}},\bm{F}_{m}^{\text{sem}}\}$ could allow for the reconstruction of the original anchor image. Accordingly, all possible combinations are considered for image reconstruction. Using the combination $(\bm{F}_{a}^{\text{sty}},\bm{F}_{m}^{\text{sem}})$ as an example, the fusion process is formulated as follows:
@@ 2.1.4. Training
During training, we enforce the model to reconstruct the original image using any combination from $\mathcal{Z}=\{\bm{z}{(a,a)},\bm{z}{(a,m)},\bm{z}{(s,a)},\bm{z}{(s,m)}\}$. This encourages the StyleTower to extract the shared information—explicitly defined as style during the construction of contrastive samples—between the anchor and positive-style images, and the SemTower to capture the shared semantics between the anchor and positive-semantic images, thereby promoting meaningful disentanglement.
@@ 2.2. Personalized Modeling
Equipped with the well-trained disentanglement mechanism, DRC facilitates personalized image generation by constructing user-specific latent instructions from the historically interacted images and the given reference image to steer generation using the LMM, as shown in Figure 3 (b). The history images provide user-preferred style information, while the reference image supplies users’ semantic intentions. Next, we describe how user-specific latent instructions are constructed for personalized image generation, followed by the design of the training and inference procedures.
@@ 2.2.1. User-specific Latent Instructions
where $\bm{m}_{i}^{\text{sty}}$ and $\bm{m}^{\text{sem}}$ denote the learned masks for $\bm{F}_{i}^{\text{sty}}$ and $\bm{F}^{\text{sem}}$. Following (Xu et al., 2024c), the masks are generated using a transformer encoder-based mask generator that predicts binary masks. Formally,
@@
$\bullet$ Obtaining Latent Instructions via Style and Semantic Fusion. Finally, we fuse the obtained user-preferred style representation with the reference semantic representation to form the user-specific latent instruction $\tilde{\bm{z}}$ for the user $u$:
@@ 3. Experiments
DM-based TI 18.67 40.90 36.58 0.7654 7.63 0.4883 32.91 53.67 48.50 105.48
LLM-based PMG 19.16 47.34 39.15 0.7383 13.94 0.4704 18.31 45.45 37.80 84.91
LLaVA* 17.88 47.26 42.59 0.7575 8.11 0.4399 17.54 42.65 39.25 93.23
LLaVA 18.72 37.44 33.19 0.7552 7.64 0.4991 27.02 49.15 43.88 95.19
LaVIT* 18.77 53.63 50.96 0.6855 9.94 0.4732 15.49 40.76 39.09 107.53
LaVIT 16.39 40.56 40.84 0.7377 8.59 0.5919 25.74 70.80 69.93 83.39
Pigeon 22.03 61.64 57.26 0.6800 6.99 0.6531 25.74 50.66 48.34 93.60
LMM-based DRC 23.19 65.10 60.00 0.6770 6.52 0.6820 21.88 48.30 45.79 92.31
Recon 16.30 40.60 40.76 0.7370 8.84 0.4715 25.84 71.09 70.14 83.57
Reference Grd 16.93 45.00 43.71 0.6443 8.10 0.5167 28.95 100.00 100.00 -
@@
Methods CS $\uparrow$ CIS $\uparrow$ DIS $\uparrow$ LPIPS $\downarrow$ StyleSim $\downarrow$ CSD $\uparrow$ CS $\uparrow$ CIS $\uparrow$ DIS $\uparrow$ FID $\downarrow$
DM-based TI 12.41 28.29 19.18 0.7721 10.83 0.3042 33.84 43.53 39.81 79.77
LLM-based PMG 13.61 25.11 22.73 0.7692 16.69 0.3141 15.60 27.29 25.15 77.25
LLaVA* 12.24 29.60 19.73 0.7607 10.02 0.3152 14.55 31.76 21.99 73.77
LLaVA 12.62 30.64 19.33 0.7690 10.67 0.3049 30.53 48.50 41.45 54.55
LaVIT* 12.64 28.23 17.50 0.7546 8.87 0.3603 19.39 36.93 37.71 50.08
LaVIT 13.86 30.49 19.95 0.7548 10.97 0.3352 25.15 46.02 60.07 33.53
Pigeon 15.19 37.42 17.70 0.7496 10.54 0.3473 27.30 47.79 41.35 41.35
LMM-based DRC 15.24 37.53 19.26 0.7538 9.57 0.3607 25.71 48.26 43.45 42.20
Recon 13.85 30.33 19.95 0.7548 9.82 0.3589 25.29 46.08 60.52 33.74
Reference Grd 10.94 51.34 20.75 0.7502 9.89 0.3769 31.81 100.00 100.00 -
@@ 3.1.2. Baselines
We compare DRC with several DM-based, LLM-based, and LMM-based methods: 1) Textual Inversion (TI) (Gal et al., 2023) learns user preferences using a word embedding to guide image generation. 2) PMG (Shen et al., 2024) converts images into keywords to capture user preferences and intentions for guiding image generation. 3) LLaVA (Liu et al., 2023) is a representative LMM for image understanding, which can be adapted for image generation with an external text-to-image generator. 4) LaVIT (Jin et al., 2024) is another LMM that quantizes images into discrete visual tokens for both image understanding and generation. 5) Pigeon (Xu et al., 2024c), built upon LaVIT, introduces a dedicated mask generator to discard noises from history images for personalized image generation. For fair comparison, both DRC and Pigeon are evaluated under the supervised fine-tuning setting only.
@@ 3.1.3. Evaluation Metrics
  • Style Alignment. To assess style preservation, we measure the visual similarity between generated images and user history images. Specifically, LPIPS (Zhang et al., 2018) is used to evaluate perceptual similarity, while StyleSim (Huang and Belongie, 2017) and CSD (Somepalli et al., 2024) are used for style-level similarity. We further compute the CLIP Image Score (CIS) and DINO Image Score (DIS) by measuring the cosine similarity of visual features extracted by CLIP (Radford et al., 2021) and DINO (Oquab et al., 2023), respectively. In addition, the CLIP Score (CS) is calculated between the generated images and textual descriptions of the history images.
@@ 3.1.4. Implementation Details
All baselines are trained using their default settings. For DRC, we set the learning rates to $1e^{-4}$ and $1e^{-5}$ during the disentanglement learning and personalized modeling stages, respectively. Following (Xu et al., 2024c), the style mask ratio $\alpha_{s}$ is fixed at 0.20.20.2. During inference, we select the optimal semantic mask ratio $\alpha_{m}\in\{0.0,0.1,\dots,1.0\}$ for each image by averaging the style CIS and the semantic CS. All experiments are conducted on a single NVIDIA A100 GPU. Detailed computational costs of DRC are reported in Appendix A.3.
@@ 3.2. Overall Performance (RQ1)
We present the performance comparison between DRC and all baselines in Table 2 with the observations as follows:
  • DM-based TI outperforms other baselines in semantic alignment by effectively capturing the semantics of the reference image, while it struggles to extract user preferences from the history images, leading to unsatisfactory style alignment.
@@
  • For the pre-trained LLaVA and LaVIT, it is challenging to effectively extract task-relevant information from complex history images and multimodal instructions, resulting in unsatisfactory performance. After fine-tuning, these models tend to focus on reconstructing the reference image while overlooking the style preferences in the user-interacted history images, which can be attributed to the lack of high-quality supervised datasets.
  • DRC and Pigeon achieve the best overall performance, excelling in both style and semantic alignment. Notably, DRC achieves comparable, even superior performance to Pigeon in most metrics. The explicit disentanglement strategy allows the model to effectively separate and integrate style and semantic information, resulting in substantial performance gains. More importantly, as illustrated in Figure 4, DRC can accurately extract and fuse user-preferred styles and reference semantics for personalized sticker generation, effectively mitigating the guidance collapse issue observed in Pigeon. For instance, in the first case, DRC successfully captures the owl character style from the user-interacted history images and combines it with the shopping semantics from the reference image to produce a shopping-themed owl sticker. In contrast, Pigeon leans toward reconstructing the reference image, neglecting the personalized style features present in the history. Additional qualitative examples are provided in Appendix A.2. Furthermore, we conduct a human evaluation to validate the effectiveness of DRC, with results presented in Appendix A.4.
@@ 3.3.1. Effect of disentanglement learning (RQ2)
Table 3. Effects of the designs in disentanglement learning. “Imp.” denotes importance sampling, “Fusion” refers to the fusion network, and“Attn.” denotes the attention-based disentangler.
@@ 4. Related Work
$\bullet$ Personalized Generation. With the rapid development of large generative models, such as DMs (Podell et al., 2024; Esser et al., 2024), LLMs (Grattafiori et al., 2024; Guo et al., 2025; Yang et al., 2024a), and LMMs (Zhou et al., 2025; Chen et al., 2025; Xie et al., 2025), personalized generation has attracted growing attention across diverse modalities, including text (Salemi et al., 2024; Qiu et al., 2025), image (Kang et al., 2025; Li et al., 2025; Xiao et al., 2025; Wu et al., 2025), 3D (Raj et al., 2023; Liu et al., 2024), video (Wang et al., 2024a; Wei et al., 2024b; Wu et al., 2024), audio (Li et al., 2024; Plitsis et al., 2024), and cross-modal generation (Pham et al., 2024; Nguyen et al., 2024). By synthesizing content that aligns with specific user preferences and requirements, personalized generation paves the way for more engaging and user-centric experiences across various application scenarios (Xu et al., 2025).
@@
$\bullet$ Large Multimodal Models. Recently, an increasing number of studies have explored LMMs for visual understanding and generation. Some approaches, such as LLaVA (Liu et al., 2023) and Mini-GPT4 (Zhu et al., 2024), leverage pre-trained vision encoders to extract visual features, which are then projected into LLMs via a dedicated mapping layer for image understanding. To enable both visual understanding and generation, models like LaVIT (Jin et al., 2024) and SEED-X (Ge et al., 2024) introduce dedicated visual tokenizers that convert images into discrete visual tokens through a learnable codebook, supporting autoregressive generation of both textual and visual tokens. In contrast, DreamLLM (Dong et al., 2024) enhances LLMs by introducing a special token to indicate image generation positions, along with a series of queries that encode historical semantics to guide the image generation process within a diffusion model-based decoder. More recently, advanced LMMs with native image generation capabilities, such as Chameleon (Team, 2024), Show-o (Xie et al., 2025), Transfusion (Zhou et al., 2025), GPT-4o (OpenAI, 2024), and BAGEL (Deng et al., 2025), have demonstrated impressive performance in multimodal understanding and generation, establishing a solid foundation for future advancements in personalized image generation. Notably, our proposed DRC adopts LaVIT as its backbone, while its design remains compatible with other LMM architectures, which we leave as an exciting direction for future exploration.